OpenStudy (anonymous):

Find the functional values f(-9), f(1), and f(8) for the compound function. f(x)={|x-5|, if x<=1 f(x)={1/x+1, if x>1 plz help, tutoring this kid and need the answer and work

jimthompson5910 (jim_thompson5910):

f(-9) means you're finding the value of f(x) when x = -9

jimthompson5910 (jim_thompson5910):

the rules state that f(x) = |x-5| if x <= 1 so this means f(x) = |x-5| f(-9) = |-9-5| f(-9) = ???

jimthompson5910 (jim_thompson5910):

f(1) is found using the same function

jimthompson5910 (jim_thompson5910):

f(8) is found using 1/x+1 because this function is used when x > 1

OpenStudy (anonymous):

@jim_thompson5910 f(-9)=-14

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

but the absolute value is never negative

jimthompson5910 (jim_thompson5910):

f(-9) = |-9-5| f(-9) = |-14| f(-9) = 14

OpenStudy (anonymous):

o mind if i ask y?

jimthompson5910 (jim_thompson5910):

absolute value is basically the distance the number is from 0

jimthompson5910 (jim_thompson5910):

-14 is 14 units away from 0 on the number line, so that's why |-14| = 14

jimthompson5910 (jim_thompson5910):

the number 76 is 76 units from 0, so |76| = 76 etc etc

OpenStudy (anonymous):

oh so f(1)=|-4| f(1)=4

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

awesome ty

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

@jim_thompson5910 f(8)=9/8 f(8)=1 1/8

jimthompson5910 (jim_thompson5910):

is 1/x+1 really \[\large \frac{1}{x}+1\] or \[\large \frac{1}{x+1}\] ?

OpenStudy (anonymous):

the 2nd one

jimthompson5910 (jim_thompson5910):

\[\large f(x) = \frac{1}{x+1}\] \[\large f(8) = \frac{1}{8+1}\] \[\large f(8) = \frac{1}{9}\]

OpenStudy (anonymous):

my bad simple error, it's been a long day lol thanks!!!!!

jimthompson5910 (jim_thompson5910):

that's ok, you're welcome